OpenStudy (anonymous):

Graph the line passing through the given point and having the given slope. Give the slope-intercept form of the equation of the line if possible. (1,-1),m = -3/2 I know that the formula is y=mx+b to solve this. And I think that the correct way of setting it up is -1=-3/2(1)+b so, when I work the problem doing this I get the following: -1=-3/2(1)+b which gives me the answer of b=1/2. After doing this I then work the problem like this: y=(-3/2)x+(1/2) replacing the y with 0 to find the x, then replacing the x with 0 to find the y. In doing so, I get that y= 1/2 and x=1/3. My problem is now that I am going through a school online, and the graph they have in the online workbook does not have the 1/2 or 1/3 on the graph. You can only do whole numbers on the graph, not fractional ones. The first point I am suppose to plot is (1,-1) then use the slope to find the 2nd one.

OpenStudy (anonymous):

you are on the right path, what you are doing is trying to find b so that you can plug it into the formula y = mx + b

OpenStudy (anonymous):

If I work that out, I get 1/2.

OpenStudy (anonymous):

so solve for b and then plug that into y = (-3/2)x + b

OpenStudy (anonymous):

and that is the slope intercept equation for that line

OpenStudy (anonymous):

How would I graph it with a slope-intercept of 1/2?

OpenStudy (anonymous):

a slope-intercept of 1/2 does not make sense. slope intercept is the form y = mx + b

OpenStudy (anonymous):

you found that b = 1/2

OpenStudy (anonymous):

Yes, but I also have a graph beside this problem that I am suppose to use to graph this.

OpenStudy (anonymous):

you need to graph y = (-3/2)x + (1/2) to do this you need to use this equation to find two points, and then draw a line through them

OpenStudy (anonymous):

two easy points to find are the intercepts, so the point where x = 0, and the point where y = 0

OpenStudy (anonymous):

So I would do 0=(-3/2)x+(1/2)

OpenStudy (anonymous):

yes you have set y = 0 which will find the x intercept of this line which will look like (some number, 0). That is one point. To find the other, set x = 0 and then solve for y to find the y intercept which will look like (0, some other number)

OpenStudy (anonymous):

ok so you have (1/3 , 0 ) and (0, 1/2) graph these two points and draw a line through them

OpenStudy (anonymous):

My graphing thing won't do 1/2 or 1/3 only whole numbers.

OpenStudy (anonymous):

are you using graph paper?

OpenStudy (anonymous):

I do it through an online school.

OpenStudy (anonymous):

what you can do in this case is try using whole number values of x to find whole number values of y, just throuhg trial and error, for example if you use x = 1, y = -1, so if you can find another point you should be able to graph it

OpenStudy (anonymous):

If I do 2=(-3/2)x+(1/2) I get that x=-1. Then I do y=(-3/2)-1+(1/2) and get that y=-2. Is this right?

OpenStudy (anonymous):

the first one is right, but youve made a mistake for your second point, (-3/2)(-1) should be a positive number

OpenStudy (anonymous):

Ok, cause it said to use the 1,-1 as the first point then find the second point using the slope.

OpenStudy (anonymous):

After I worked all this out.

OpenStudy (anonymous):

so you just need to find one more point?

OpenStudy (anonymous):

i mean the online thingy is just asking for another x and y value?

OpenStudy (anonymous):

What it says is the very question I posted word for word, but after I work out two points then submit it, it says it's wrong that I need to use (1,-1) as one of the points.

OpenStudy (anonymous):

what are the two points you are submitting?

OpenStudy (anonymous):

The two I did up above, but I made the second one into a positive.

OpenStudy (anonymous):

you've entered (1,-1) and (-1,2) and it says its wrong?

OpenStudy (anonymous):

There it went through this time, probably a mistake I was doing with the signs. Now I just have to write the slope intercept equation of the line.

OpenStudy (anonymous):

slope intercept is the form y = mx +b but you'll plug in the values for m and b

OpenStudy (anonymous):

So it would be -1=-3/2x+1/2

OpenStudy (anonymous):

when writing the equation of a line in slope intercept form, you leave x and y as variables

OpenStudy (anonymous):

the form you have written only describes a specific point on the line, because you've plugged in a value for y, to represent the entire lin e you leave x and y as variables